### v1

# Structure Guided Large Language Model for SQL Generation

Abstract
Generating accurate Structured Querying Language (SQL) is a long-standing problem, especially in matching users’ semantic queries with structured databases and then generating structured SQL. Existing models typically input queries and database schemas into the LLM and rely on the LLM to perform semantic-structure matching and generate structured SQL. However, such solutions overlook the structural information within user queries and databases, which can be utilized to enhance the generation of structured SQL. This oversight can lead to inaccurate or unexecutable SQL generation. To fully exploit the structure, we propose a structure-to-SQL framework, which leverages the inherent structure information to improve the SQL generation of LLMs. Specifically, we introduce our Structure Guided SQL (SGU-SQL) generation model. SGU-SQL first links user queries and databases in a structure-enhanced manner. It then decomposes complicated linked structures with grammar trees to guide the LLM to generate the SQL step by step. Extensive experiments on two benchmark datasets illustrate that SGU-SQL can outperform sixteen SQL generation baselines.
[…]
Structure Guided Large Language Model for SQL Generation
Qinggang Zhang1, Junnan Dong1, Hao Chen1, Wentao Li1, Feiran Huang2, Xiao Huang1thanks:   Corresponding author The Hong Kong Polytechnic University, 2Jinan University (qinggangg.zhang,hanson.dong,wentao0406.li)@connect.polyu.hk sundaychenhao@gmail.com; huangfr@jnu.edu.cn; xiaohuang@comp.polyu.edu.hk
1 Introduction
Structured Querying Language (SQL) is the most popular language for communicating with databases and retrieving user-desired data. Due to the popularity of Large Language Models (LLMs), there is an increasing need to input a user query and automatically generate the corresponding SQL to communicate with the database. However, unlike natural languages, SQL is a standardized structured language that must exactly match the database in order to be executable and retrieve the correct data. Considering the impact of these factors, incorporating structure information can improve the performance of SQL generation models.
Existing state-of-the-art solutions primarily rely on the comprehensive and generative capabilities of LLMs to generate SQL, thus adopting a text-to-SQL approach. T5 Raffel et al. (2020) first attempts to train a language model to understand the user query and build connections with the database. After the emergence of ChatGPT, especially GPT4 OpenAI (2023), some pioneers Dong et al. (2023b); Ma et al. (2023) directly input the user query and database schema to ChatGPT to generate SQL. Building upon this, DIN-SQL Pourreza and Rafiei (2023) employs chain-of-thought by requiring ChatGPT to decompose the SQL generation into several sub-tasks and then lead ChatGPT to generate the SQL according to the sub-tasks.
However, existing models primarily rely on the semantic information of user queries and database schema, while overlooking the inherent structure of queries, databases, and SQLs. The neglect of structure results in the following limitations:
1. Semantic Ambiguity: Only relying on the semantics of the query and the database may not fully capture the user’s intent. For example, when a user queries “income”, it may ambiguously refer to the summation of the “salary” and “bonus” tables in the database. However, these tables are indirectly linked through the "ID" and "DT" (DateTime) keys. By adopting a purely semantic approach, there is a risk of overlooking this implicit linking and neglecting either the salary or the bonus.
2. Complex Relational Linking: Real-world databases often contain implicit and intricate structures, such as storing metadata in meta tables and storing transactional data in detail tables. The complicated structure presents challenges for semantic-based SQL generation in precisely organizing the complex structure.
3. Opaque Matching Processes: Existing models employ a black-box approach to query-database matching, which provides little interpretability to users. As a result, it is difficult for users to inspect or refine the SQL generation process, thus limiting the reliability and usability.
While leveraging structural information to enhance SQL generation shows promise, it remains a challenging task due to the following reasons.
1. Structured Representation: The challenge lies in creating representations that accurately capture the structural nuances of queries and databases. It is especially demanding to develop a model that aligns query structures with database architectures.
2. Precise Query Matching: The key to SQL generation is linking query intents precisely with database structures. It is challenging to propose an accurate structured query-database linking.
3. Decomposition Strategies: When dealing with complex queries, it is challenging to devise effective strategies for decomposing the complicated linked structure for step-by-step SQL generation.
In this paper, we enhance the existing framework of text-to-SQL by introducing the tailored Structure Guided SQL (SGU-SQL) generation model. This model represents user queries and databases in a structured manner and employs a tailored semantic-enhanced structure matching model to establish a connection between the query structure and the database structure. Subsequently, we decompose the linked structure into sub-grammar trees to guide the step-by-step generation of SQLs by the LLM. The main contributions of our research are summarized as follows:
  • We discover the limitation of neglecting the structure in the SQL generation task and propose our tailored Structure Guide to utilize the structure information to enhance the SQL generation.
  • SGU-SQL proposes graph-based structure construction to comprehend user query and database understanding and then link query structure and database structure with semantic-enhanced solutions.
  • SGU-SQL introduces tailored structure-decomposed generation strategies to decompose queries with grammar trees and then generate accurate SQL with LLM.
  • Extensive experiments on two benchmark datasets verify SGU-SQL outperforms sixteen state-of-the-art SQL-generation baselines.
2 Related Work
Text-to-SQL facilitates user interaction with databases via natural language queries, circumventing the necessity for SQL syntax and database structure proficiency. Historically, this challenge was tackled through rule-based methods and hand-crafted heuristics. However, these approaches often faltered due to the complexity and diversity inherent in natural language queries.
Figure 1: The overall framework.
2.1 Fine-tuning Large Language Models for Text-to-SQL
The emergence of large language models (LLMs) has markedly altered the landscape for text-to-SQL tasks. LLMs, with their capacity for understanding and generating human-like text, present a robust solution for text-to-SQL applications Liu et al. (2023). The development of LLMs typically encompasses pre-training followed by fine-tuning. Research has concentrated on fine-tuning with domain-specific data and optimization techniques to enhance base models for coding tasks, including text-to-SQL. This process enables models to master programming language syntax and database schema intricacies Raffel et al. (2020); Roziere et al. (2023). Through training on tailored datasets of annotated SQL queries, LLMs acquire the grammar and structure necessary for generating compliant SQL code Trummer (2022); Sun et al. (2023). Furthermore, Scholak et al. (2021) introduced a decoding mechanism for LLMs that ensures the generation of valid sequences by discarding inadmissible tokens at each step, employing incremental parsing to guarantee the validity of SQL queries produced by auto-regressive models.
2.2 In-Context Learning for Text-to-SQL
In-context learning enhances LLM performance by providing contextual examples during inference, thereby improving performance for specific tasks Zhang et al. (2023a). This approach has seen innovative applications in text-to-SQL, with strategies aimed at optimizing example selection based on user queries and referenced data. Liu et al. (2021) first investigated the impact of in-context examples on GPT-3’s performance. Nan et al. (2023) proposed categorizing problems by complexity to facilitate the identification of similarly complex examples, whereas Guo et al. (2023) developed a domain-specific vocabulary masking technique for example similarity assessment, highlighting the relevance of SQL-specific terms. DIN-SQL represents a sophisticated application of this concept, dissecting the text-to-SQL task into schema linking, query classification, and decomposition to improve the generation of complex SQL queries Pourreza and Rafiei (2023). DAIL-SQL further enhanced the performance by incorporating suitable formatting of the database schema and selecting examples based on skeleton similaritiesGao et al. (2023). Dong et al. (2023b) designed a framework for ChatGPT with clear prompting for effective input layout and tailored hints for calibration and consistency check during the query generation.
2.3 Structure Learning for Text-to-SQL
As text-to-SQL models often employ an encoder-decoder architecture, many strategies such as Graph Neural Networks (GNNs) Chen et al. (2022); Zhang et al. (2023b, 2022); Dong et al. (2023a); Chen et al. (2020); Shengyuan et al. (2024) have been designed to learn structural representations for the decoders to comprehend questions and database schemas. RATSQL employs a graph-based structure to delineate relationships within database schemas and queries, treating the schema as a graph of tables and columns connected by relational edges Wang et al. (2019). Cao et al. (2021) introduced LGESQL, an edge-centric graph model derived from conventional node-centric graphs, to capture diverse structural topologies. S2SQL, proposed by Hui et al. (2022), integrates syntactic dependency information into a question-schema interaction graph, focusing on primary relationships to mitigate overfitting while emphasizing essential graph structures. Li et al. (2023b) explored the integration of GNN layers into the T5 model, aiming to leverage both semantic and structural information from PLMs and GNNs, respectively. Li et al. (2023a) designed a ranking-enhanced encoder to rank and filter the schema items for the skeleton-aware decoder to decouple the schema linking and the skeleton parsing stage.
3 Methodology
In this section, we introduce the rationales of our framework. We leverage the implicit structural information in both queries and databases from three aspects: $(i)$ A graph-based structure construction for both user query and database understanding; $(ii)$ A tailored structure linking method is proposed to bind the query nodes with corresponding schema nodes, i.e., tables and columns. $(iii)$ Structure prompting to LLMs for accurate SQL generation. The overall illustration is presented in Figure 1.
3.1 Structure Construction
Initiating LLMs with the comprehension of the user query and database schema is challenging in complex real-world scenarios. First, natural language queries often contain ambiguous references or terminologies. Second, the implicit structure information is not readily accessible in the plain text format. In this part, we introduce the details about structure extraction in both query and database to achieve effective mapping and interaction.
Inspired by the superiority of graphs, which can effectively represent the structural information in both natural language sentences and real-world entities Dong et al. (2023a), we aim to construct build a comprehensive query-schema graph by combining the query structure and databse structure based on the topic concepts in the query, pre-defined relations in the query, database schema as well as the table/column names present within the schema. The graph contains three key structures:
  • Query Structure ($R_{\text{query}}$): Encodes syntactic dependencies between tokens in the question, derived from its syntactic parse.
  • Database Structure ($R_{\text{schema}}$): Represents intrinsic relationships within the database schema, like foreign keys.
  • Linking Structure ($R_{\text{linking}}$): Aligns query entities with corresponding schema columns or tables.
To make it clear, we introduce the implementation details of query and database structure construction in Section 6.1 of the Appendix, and propose a novel structure learning framework to obtain the linking structure in the following section.
3.2 Structure Linking
Schema linking plays a vital role in accurately interpreting and generating SQL queries from textual input.Existing schema linking has been actively adopted as a rule-based mapping problem. However, the hard mapping based on concept and table/column names results in an unsatisfactory performance. To tackle this, given the constructed query and database structures, we value the adjacency information during the matching process and propose to automatically build the connection between the query structure and schema at the node level. Specifically, we design a tailored structure-based linking framework. Both query and schema structures are first encoded through a Relational Graph Attention Transformer (RGAT) for initial node representations. The representation learning process is guided by the message propagation within the self-structure. We formalize the procedure of structure-aware question-schema structure linking as follows:
$\mathcal{G}_{s}^{Agg}={\rm Agg}(\mathcal{G}_{s},\mathcal{G}_{q}),$ (1)
$\mathcal{G}_{q}^{Agg}={\rm Agg}(\mathcal{G}_{q},\mathcal{G}_{s}),$ (2)
where the structure-aware aggregation function ${\rm Agg}(.)$ is employed to gather information from both the schema-graph $\mathcal{G}_{s}$ and the query-graph $\mathcal{G}_{q}$ and transfer it to the adjacent graph.
Let $\{\boldsymbol{h}_{i}^{q}\}^{m}_{i=1}$ represent a set of node embeddings in the query-graph $\mathcal{G}_{q}$ and let $\{\boldsymbol{h}_{j}^{k}\}^{n}_{j=1}$ denote a set of node embeddings in the subgraph $\mathcal{G}_{k}$ that extracted from schema-graph $\mathcal{G}_{s}$. In particular, we first employ global-average pooling on the node embedding $\boldsymbol{h}_{i}^{q}$ of the query structure $\mathcal{G}_{q}$ in order to derive the global query structure embedding $\boldsymbol{h}^{q}_{g}$. Following this, to encapsulate globally pertinent information, the key node embedding $\boldsymbol{h}^{k}_{j}$ is updated in the subsequent manner:
$\boldsymbol{h}^{q}_{g}=\frac{1}{m}\sum\nolimits^{m}_{i=1}\boldsymbol{h}^{q}_{i},$ (3)
$\alpha_{j}=\theta\left({\boldsymbol{h}^{q}_{g}}^{T}\boldsymbol{W}_{g}\boldsymbol{h}^{k}_{j}\right),$ (4)
$\boldsymbol{h}^{k}_{j}=\sum\nolimits_{l\in\mathcal{N}_{j}}\alpha_{l}\boldsymbol{W}_{k}\boldsymbol{h}^{k}_{l}+\alpha_{j}\boldsymbol{W}_{k}\boldsymbol{h}^{k}_{j}$ (5)
$+(1-\alpha_{j})\boldsymbol{W}_{q}\boldsymbol{h}^{q}_{g}$ (6)
where $\boldsymbol{W}_{g}$, $\boldsymbol{W}_{q}$, $\boldsymbol{W}_{k}$ represent trainable parameters, and $\theta$ illustrates a sigmoid function. While $\alpha_{j}$ denotes the relevance score situated between the $j$-th key node and the global query structure.
For each node $a$ in the query structure $\mathcal{G}_{q}$, it is necessary to find a corresponding matching node $s$ in the database $\mathcal{G}_{s}$. The proposed solution mainly consists of three steps. First, a set of most relevant candidate nodes $\{s_{1},s_{2},\dots,s_{K}\}$ is identified through string matching in the set of tables and columns $A$. Second, for each candidate node $s$, a specialized subgraph $\mathcal{G}(a,s)$ is constructed. As shown in Figure 1, $\mathcal{G}(a,s)$ includes the query graph $\mathcal{G}_{q}$, adjacent nodes of $s_{k}$ (first-order and second-order), and an edge connecting $a$ and $s_{k}$. Lastly, we train a structure learning model $\text{RGAT}(\cdot)$ to learn the graph-level embedding representation of $\mathcal{G}_{(}a,s_{k})$ for predicting a matching score of the candidate pair $(a,s_{k})$ The expression for this is as follows,
$\boldsymbol{h}=\text{RGAT}(\mathcal{G}(a,s)),$ (7)
$\text{Score}_{(a,s_{k})}=\sigma(\sum\nolimits_{l\in\mathcal{G}(a,s_{k})}\boldsymbol{h}^{k}_{l}).$ (8)
Based on positive samples $(a,s)$ and negative samples $(a,s_{k})$, where $s_{k}\neq s$, the structure learning model $\text{RGAT}(\cdot)$ is iteratively trained.
Figure 2: The example of grammar tree. The subfigure $(a)$ denotes the context-free grammar of SQL. While subfigures $(b)$ and $(c)$ demonstrate two examples of the constructed grammar tree used in our SGU-SQL.
3.3 Structure-Decomposed Prompting
Decomposing query with grammar tree The performance of LLMs on complex tasks can be improved by using decomposing-based methods. However, decomposing a SQL query into subtasks is challenging due to its declarative structure and the intricate connections between query concepts. To this end, in this section, we introduce a context-free grammar tree (defined in Figure 2-(a)) to break the query into subtasks according to the syntax structure of the user query. Specifically, we first employ the Stanford Parser Klein and Manning (2003) to build the grammar tree to achieve a linguistic understanding of the natural language query and then adopt a node mapper to match nodes in the linguistic parse tree to SQL operations Kate (2008). Following this, the original query can be divided into several subtasks according to the SQL operations distributed on the grammar tree. More details are presented in Section 6.3 of Appendix. Meta-operation in SQL Based on the commonly used operations and properties in SQL, we define four meta-operations as the major identifiers for sub-task division:
  • A: a start of a SELECT statement;
  • R: a root node or continue with the result returned by last SELECT;
  • T: a corresponding Table name;
  • C: a corresponding Column name.
Following this, the original query can be divided into several subtasks according to the SQL operations distributed on the grammar tree.
Prompt design for Text-to-SQL Given LLMs are sensitive to the prompts, we aim to transform the graph-structured sub-tasks into the form of text. This helps the LLMs better understand the task requirements to generate effective SQL queries. Specifically, given a user query $q_{j}$, to systematically guide the language model in generating SQL queries, we decompose the generation task into multiple steps and guide the language model to solve it step by step. The final prompts follow the format $<Q_{j};G_{j};<Q_{j1};A_{j1}\dots;Q_{jk};A_{jk}>;I_{j};A_{j}>$, where $k$ represents the number of subtasks, $Q_{j}$ and $A_{j}$ represent the statement of each subtask and the relevant information involved in this task, respectively, $G_{j}$ denotes the linked query-schema structure, and $I_{j}$ is the task description.
4 Experiments
This section empirically evaluates the proposed framework, i.e., SGU-SQL, and presents its performance on two benchmark datasets. Our empirical study is motivated by the following questions:
  • Q1 How does our proposed SGU-SQL perform in comparison with the strongest baselines, including traditional finetuning-based, structure-learning-based methods, and other in-context-learning-based methods?
  • Q2 Could our proposed SGU-SQL enhance other LLMs by substituting the original framework with the structure-decomposing-based prompt?
  • Q3 Is our proposed structure prompting effective when handling queries of different difficulty levels? What types of queries are prone to errors in our model? What is the reason for the error?
4.1 Experiment Setup
Table 1: The Execution Accuracy and Exact Match Accuracy of text-to-SQL models on Spider and Bird.
Model Spider Bird
Exec Acc EM Acc Exec Acc EM Acc
Fine-tuning-based
T5-Base Raffel et al. (2020) 0.5791 0.5722 0.1154 0.1009
T5-Large Raffel et al. (2020) 0.6728 0.6536 0.1970 0.1736
T5-3B Raffel et al. (2020) 0.7440 0.7152 0.2334 0.2068
Structure-learning-based
RAT-SQL + GAP Wang et al. (2019) 0.7505 0.7379 0.2639 0.2431
RASAT + PICARD Qi et al. (2022) 0.8048 0.7537 0.2870 0.2603
[…]
RESDSQL-3B Li et al. (2023a) 0.8173 0.7580 0.3031 0.2569
Graphix-3B + PICARD Li et al. (2023b) 0.8121 0.7563 0.2984 0.2418
Incontext-learning-based
CodeX Chen et al. (2021) 0.7167 0.4905 0.3438 0.3019
PaLM-2 Anil et al. (2023) 0.6985 0.4438 0.2735 0.2543
ChatGPT OpenAI (2023) 0.7394 0.5327 0.3562 0.3041
GPT-4 OpenAI (2023) 0.7665 0.5892 0.4633 0.4255
CoT + ChatGPT Wei et al. (2022) 0.7552 0.5684 0.3665 0.3127
C3-GPT Dong et al. (2023b) 0.8185 0.7036 - -
DIN-SQL + GPT-4 Pourreza and Rafiei (2023) 0.8237 0.7187 0.5079 0.4398
DAIL-SQL + GPT-4 Gao et al. (2023) 0.8449 0.7443 0.5434 0.4581
SGU-SQL + GPT-4 (Ours) 0.8791 0.7679 0.5771 0.4989
Datasets We assess the performance of text-to-SQL models using two renowned datasets, Spider Yu et al. (2019) and BIRD Li et al. (2023c). Spider, a cross-domain text-to-SQL dataset, comprises 8659 instances in the training split and 1034 instances in the development split, spanning across 200 databases. Each instance comprises a natural language question related to a specific database and its corresponding SQL query. For evaluation purposes, we utilize the Spider-dev development split since the test split has not been released. On the other hand, BIRD (BIg Bench for large-scale Database Grounded text-to-SQL Evaluation) is another pioneering cross-domain dataset that focuses on exploring the impact of extensive database contents on text-to-SQL parsing. BIRD features over 12,751 unique question-SQL pairs, encompassing 95 large databases with a total size of 33.4 GB. It encompasses more than 37 professional domains, including blockchain, hockey, healthcare, education, and others.
[…]
Evaluation Metrics The effectiveness of our models is analyzed by employing two frequently used measurement standards, namely Exact-Set-Match Accuracy (EM Acc) and Execution Accuracy (Exec Acc). EM Acc approaches treat each clause as an individual set and examine its prediction to the equivalent validated SQL query. However, this parameter does not take into consideration values, potentially triggering false positives and negatives. Conversely, Exec Acc places the execution result of the prospective SQL query alongside that of the confirmed SQL query on certain database iterations. With Execution Accuracy, a more nuanced understanding of the model’s functionality is attained since there can be several legitimate SQL queries corresponding to one question, and Exact-Set-Match Accuracy only assesses the predicted SQL in contrast to one of them.
Execution accuracy Model Easy Medium Hard Extra All SGU-SQL + GPT-4 93.5 91.9 845 72.3 87.9 DAIL-SQL + GPT-4 90.6 82.7 75.4 58.2 84.5 DIN-SQL + GPT-4 91.5 78.4 69.4 54.3 82.4 GPT-4 87.3 76.0 63.2 51.9 76.7 CodeX 82.8 67.9 57.7 48.5 73.9 Exact match accuracy Model Easy Medium Hard Extra All SGU-SQL + GPT-4 85.5 83.9 72.69 51.2 76.8 DAIL-SQL + GPT-4 79.5 74.2 65.2 46.2 74.4 DIN-SQL + GPT-4 81.3 67.1 42.7 46.1 71.9 GPT-4 77.9 56.4 37.9 22.5 58.9 CodeX 65.7 42.5 28.7 15.1 49.1
Table 2: Performance of SGU-SQL and state-of-the-art in-context learning competitors on the development set of Spider with queries of different difficulty levels.
4.2 Main Results: Q1
To answer Q1, we perform extensive experiments to compare SGU-SQL with the strongest baselines best to our knowledge. The comparison results on Spider and Bird are placed in Table 1. We summarize our experimental observations as follows.
Obs.1. Our proposed SGU-SQL significantly outperforms SOTA text-to-SQL baselines. We compare SGU-SQL (ours) with $3$ fine-tuning based methods, $5$ structure-learning-based methods and $8$ SOTA incontext-learning-based methods over two benchmark datasets. As shown in Table 1, SGU-SQL achieves superior performance across both datasets, outperforming all baselines in terms of execution accuracy and exact match accuracy.
Obs.2. In-context learning-based method is better than the methods of the other two categories. Among the three categories of methods, in-context learning-based methods consistently demonstrate superior performance. This suggests that leveraging in-context learning mechanisms is crucial for enhancing the understanding and generation of SQL queries from natural language inputs. Specifically, the in-context learning-based methods, i.e., DIN-SQL and DAIL-SQL in our comparison set achieve higher accuracy rates and require less computational overhead compared to fine-tuning and structure-learning-based methods. Additionally, the in-context learning-based methods exhibit better generalization across different datasets, indicating their robustness and adaptability.
4.3 Ablation Staudy: Q2
The effect of prompting strategy In this part, we conduct comprehensive experiments to investigate the effectiveness of our proposed prompting strategy. Specifically, we compare the structure-based decomposing strategy used in our SGU-SQL with other prompting strategies like CoT Wei et al. (2022) and few-shot prompting. As shown in Table 3, we can have the following observations.
Obs.1. Our structure-based decomposing significantly outperforms other prompting strategies. Our method demonstrates superior performance across all tested LLMs. Specifically, compared to CoT, our approach achieves an average improvement of 5.03%, while outperforming few-shot prompting by 4.98% on average.
Obs.2. Traditional decomposing-based methods are ineffective in the text-to-SQL task. While decomposing complex tasks into subtasks like CoT, can enhance model performance in many natural language understanding tasks, it proves to be ineffective in the text-to-SQL task. As shown in Table 3, applying COT on PaLM-2 even leads to a performance decrease of 1.08% compared to the naive few-shot prompting. This is attributed to the complex syntax of SQL, and the intricate correspondence between query terms in user queries and database data units. Conversely, we formally define the meta-operations in SQL and propose a decomposing strategy according to the grammar tree to separate the query into subtasks. This boosts the LLMs’ comprehension of linked queries to generate accurate SQLs step by step.
The generalization ability of prompts To further verify the generalization ability of our proposed prompting strategy, in this part, we conduct comprehensive experiments to investigate whether SGU-SQL could enhance other LLMs by substituting their original framework with the decomposing-based prompts. Specifically, we replace Gpt-4 used in SGU-SQL with other representative generative LLMs, including PaLM-2 Anil et al. (2023), CodeX Chen et al. (2021), ChatGPT and GPT-4 OpenAI (2023) as alternatives. Specifically, we used the model ‘chat-bison-001’ provided by GoogleAI as the implementation of PaLM-2, and ‘gpt-3.5-turbo’ and ‘gpt-4’ as the implementations of ChatGPT and GPT-4, respectively. The text-to-SQL task is conducted under the few-shot setting with the query from the development set of Spider as input. As shown in Figure LABEL:figure2, we have the following observations.
Obs.1. The performances of the original LLMs improved significantly by integrating the prompt learned from our SGU-SQL. Specifically, PaLM-2 improved by 4%, CodeX by 3%, ChatGPT by 5%, and GPT-4 by almost 11%. The substantial performance gains indicate the robustness and generalization ability of our proposed prompting strategy. Furthermore, the consistent improvements across different LLMs highlight the versatility and applicability of our approach in enhancing the capabilities of existing language models.
Obs.2. LLMs with stronger reasoning abilities exhibit greater improvement. We observe that LLMs with stronger reasoning abilities benefit more from integrating the prompts learned from SGU-SQL. Specifically, GPT-4, which is known for its advanced reasoning capabilities, shows a more substantial performance improvement compared to PaLM-2, CodeX, and ChatGPT. This suggests that our prompting strategy is particularly effective in enhancing the performance of LLMs that require more complex reasoning tasks.
Table 3: Performance of LLMs with different prompting strategies on the development set of Spider.
LLM Prompting strategy Exec acc
PaLM-2 + Few-shot Prompting 0.6985
+ CoT Prompting 0.6873
+ SGU-SQL 0.7395
CodeX + Few-shot Prompting 0.7167
+ CoT Prompting 0.7198
+ SGU-SQL 0.7418
ChatGPT + Few-shot Prompting 0.73947
+ CoT Prompting 0.7552
+ SGU-SQL 0.7846
GPT-4 + Few-shot Prompting 0.7665
+ CoT Prompting 0.7834
+ SGU-SQL 0.8791
4.4 Model Analysis: Q3
Difficulty analysis In this part, we first analyze the performance of our proposed method on queries with different levels of difficulty. Our analysis focused on evaluating the performance of our proposed method across queries of varying difficulty levels. Table 2 provides a comparative assessment of our method against state-of-the-art (SOTA) prompting methods on the Spider development set. Our findings reveal that our method consistently outperforms competing methods across all difficulty levels. Notably, we observe the most substantial improvements in the extra hard and hard classes, where other prompting models struggle. Additionally, our method also shows slight improvement in the easy class, which suggests that our method is robust and effective across queries of different difficulty levels, highlighting its potential for practical applications in natural language understanding and query generation tasks.
Error analysis We checked the errors in the generated SQL answers and classified them into six categories, as shown in Figure LABEL:fig:error following the classification by Pourreza and Rafiei (2023). We discuss the failure cases of our model in comparison with baseline models, and then discuss the reasons for the typical failure for LLMs in Text-to-SQL tasks. Compared to the baseline model, we achieved a reduction of approximately 33.5% in errors and made progress in the schema-linking and join statement components where traditional models often falter. Please refer to Section 8 of Appendix for more detailed analysis.
5 Conclusion and Future Work
Generating accurate Structured Query Language (SQL) has long been a challenging task, especially in matching users’ semantic queries with structured databases. Existing models typically input queries and database schemas into LLMs to perform semantic-structure matching and generate structured SQL. However, these models often overlook the structural information inherent in user queries and databases, which could significantly enhance the generation of accurate SQL queries. This oversight can result in the production of inaccurate or unexecutable SQL queries. To fully exploit the structure, we propose the structure-to-SQL framework, which leverages the inherent structure information to improve the SQL generation of LLMs. Specifically, we introduce our Structure Guided SQL (SGU-SQL) generation model. SGU-SQL links user queries and databases in a structure-enhanced manner. It then decomposes complicated linked structures with grammar trees to guide the LLM to generate the SQL step by step. Extensive experiments conducted on two benchmark datasets demonstrate SGU-SQL outperforms all state-of-the-art baselines.
Limitations
The limitation of our approach is a common constraint for text-to-SQL models, as these models depend on GPT4 to generate the SQL queries. Consequently, if the network is inaccessible or the OpenAI API is blocked, the text-to-SQL models are unable to generate SQL queries as expected.
Ethics Statement
We confirm that we have fully complied with the ACL Ethics Policy in this study. We conduct experiments with widely adopted publicly available datasets. The codes and generated prompts will be open-sourced for other researchers’ fair reproduction and further study in the active text-to-SQL community.

### v2

@@
 Table 1: The Execution Accuracy and Exact Match Accuracy of text-to-SQL models on Spider and Bird.
-Model Spider Bird
-Exec Acc EM Acc Exec Acc EM Acc
-Fine-tuning-based
-T5-Base Raffel et al. (2020) 0.5791 0.5722 0.1154 0.1009
-T5-Large Raffel et al. (2020) 0.6728 0.6536 0.1970 0.1736
-T5-3B Raffel et al. (2020) 0.7440 0.7152 0.2334 0.2068
-Structure-learning-based
-RAT-SQL + GAP Wang et al. (2019) 0.7505 0.7379 0.2639 0.2431
-RASAT + PICARD Qi et al. (2022) 0.8048 0.7537 0.2870 0.2603
-[…]
-RESDSQL-3B Li et al. (2023a) 0.8173 0.7580 0.3031 0.2569
-Graphix-3B + PICARD Li et al. (2023b) 0.8121 0.7563 0.2984 0.2418
-Incontext-learning-based
-CodeX Chen et al. (2021) 0.7167 0.4905 0.3438 0.3019
-PaLM-2 Anil et al. (2023) 0.6985 0.4438 0.2735 0.2543
-ChatGPT OpenAI (2023) 0.7394 0.5327 0.3562 0.3041
-GPT-4 OpenAI (2023) 0.7665 0.5892 0.4633 0.4255
-CoT + ChatGPT Wei et al. (2022) 0.7552 0.5684 0.3665 0.3127
-C3-GPT Dong et al. (2023b) 0.8185 0.7036 - -
-DIN-SQL + GPT-4 Pourreza and Rafiei (2023) 0.8237 0.7187 0.5079 0.4398
-DAIL-SQL + GPT-4 Gao et al. (2023) 0.8449 0.7443 0.5434 0.4581
-SGU-SQL + GPT-4 (Ours) 0.8791 0.7679 0.5771 0.4989
+{adjustwidth}
+-0.5 cm-0.5 cm Model Spider Bird Exec Acc EM Acc Exec Acc EM Acc Fine-tuning-based T5-Base Raffel et al.
+(2020) 0.5791 0.5722 0.1154 0.1009 T5-Large Raffel et al.
+(2020) 0.6728 0.6536 0.1970 0.1736 T5-3B Raffel et al.
+(2020) 0.7440 0.7152 0.2334 0.2068 Structure-learning-based RAT-SQL + GAP Wang et al.
+(2019) 0.7505 0.7379 0.2639 0.2431 RASAT + PICARD Qi et al.
+(2022) 0.8048 0.7537 0.2870 0.2603 S2SQL + ELECTRA Hui et al.
+(2022) 0.7643 0.7285 - - RESDSQL-3B Li et al.
+(2023a) 0.8173 0.7580 0.3031 0.2569 Graphix-3B + PICARD Li et al.
+(2023b) 0.8121 0.7563 0.2984 0.2418 Incontext-learning-based CodeX Chen et al.
+(2021) 0.7167 0.4905 0.3438 0.3019 PaLM-2 Anil et al.
+(2023) 0.6985 0.4438 0.2735 0.2543 ChatGPT OpenAI (2023) 0.7394 0.5327 0.3562 0.3041 GPT-4 OpenAI (2023) 0.7665 0.5892 0.4633 0.4255 CoT + ChatGPT Wei et al.
+(2022) 0.7552 0.5684 0.3665 0.3127 C3-GPT Dong et al.
+(2023b) 0.8185 0.7036 - - DIN-SQL + GPT-4 Pourreza and Rafiei (2023) 0.8237 0.7187 0.5079 0.4398 DAIL-SQL + GPT-4 Gao et al.
+(2023) 0.8449 0.7443 0.5434 0.4581 SGU-SQL + GPT-4 (Ours) 0.8791 0.7679 0.5771 0.4989
 Datasets We assess the performance of text-to-SQL models using two renowned datasets, Spider Yu et al.
@@
 It encompasses more than 37 professional domains, including blockchain, hockey, healthcare, education, and others.
-[…]
+Baselines To valid the effectiveness of SGU-SQL, we compare it with the state-of-art baselines according to our best knowledge.
+Following the taxonomy aforementioned in Section 2, we divide all baselines into three categories: $(i)$ Fine-tuning-based: T5-base Raffel et al.
+(2020), T5-large Raffel et al.
+(2020), T5-3B Raffel et al.
+(2020); $(ii)$ structure-learning-based: RAT-SQL Wang et al.
+(2019), RASAT Qi et al.
+(2022), S2SQL Hui et al.
+(2022) ,RESDSQL Li et al.
+(2023a),GRAPHIX Li et al.
+(2023b); and $(iii)$ incontext-learning-based: PaLM-2 Anil et al.
+(2023), CodeX Chen et al.
+(2021), ChatGPT OpenAI (2023), GPT-4 OpenAI (2023), C3-GPT Dong et al.
+(2023b), DIN-SQL Pourreza and Rafiei (2023), DAIL-SQL Gao et al.
+(2023).
 Evaluation Metrics The effectiveness of our models is analyzed by employing two frequently used measurement standards, namely Exact-Set-Match Accuracy (EM Acc) and Execution Accuracy (Exec Acc).
@@
 With Execution Accuracy, a more nuanced understanding of the model’s functionality is attained since there can be several legitimate SQL queries corresponding to one question, and Exact-Set-Match Accuracy only assesses the predicted SQL in contrast to one of them.
-Execution accuracy Model Easy Medium Hard Extra All SGU-SQL + GPT-4 93.5 91.9 845 72.3 87.9 DAIL-SQL + GPT-4 90.6 82.7 75.4 58.2 84.5 DIN-SQL + GPT-4 91.5 78.4 69.4 54.3 82.4 GPT-4 87.3 76.0 63.2 51.9 76.7 CodeX 82.8 67.9 57.7 48.5 73.9 Exact match accuracy Model Easy Medium Hard Extra All SGU-SQL + GPT-4 85.5 83.9 72.69 51.2 76.8 DAIL-SQL + GPT-4 79.5 74.2 65.2 46.2 74.4 DIN-SQL + GPT-4 81.3 67.1 42.7 46.1 71.9 GPT-4 77.9 56.4 37.9 22.5 58.9 CodeX 65.7 42.5 28.7 15.1 49.1
+{tblr} cells = c, row1 = Silver, row8 = Silver, cell11 = c=6, cell81 = c=6, hlines, Execution accuracy &
+Model Easy Medium Hard Extra All
+SGU-SQL + GPT-4 93.5 91.9 845 72.3 87.9
+DAIL-SQL + GPT-4 90.6 82.7 75.4 58.2 84.5
+DIN-SQL + GPT-4 91.5 78.4 69.4 54.382.4
+GPT-4 87.3 76.0 63.2 51.9 76.7
+CodeX 82.8 67.9 57.7 48.5 73.9
+Exact match accuracy
+Model Easy Medium Hard Extra All
+SGU-SQL + GPT-4 85.5 83.9 72.69 51.2 76.8
+DAIL-SQL + GPT-4 79.574.265.246.274.4
+DIN-SQL + GPT-4 81.3 67.1 42.7 46.1 71.9
+GPT-4 77.9 56.4 37.9 22.5 58.9
+CodeX 65.7 42.5 28.7 15.1 49.1
 Table 2: Performance of SGU-SQL and state-of-the-art in-context learning competitors on the development set of Spider with queries of different difficulty levels.
@@
 Table 3: Performance of LLMs with different prompting strategies on the development set of Spider.
-LLM Prompting strategy Exec acc
-PaLM-2 + Few-shot Prompting 0.6985
-+ CoT Prompting 0.6873
-+ SGU-SQL 0.7395
-CodeX + Few-shot Prompting 0.7167
-+ CoT Prompting 0.7198
-+ SGU-SQL 0.7418
-ChatGPT + Few-shot Prompting 0.73947
-+ CoT Prompting 0.7552
-+ SGU-SQL 0.7846
-GPT-4 + Few-shot Prompting 0.7665
-+ CoT Prompting 0.7834
-+ SGU-SQL 0.8791
+{adjustwidth}
+-0.
+cm-0.
+cm LLM Prompting strategy Exec acc PaLM-2 + Few-shot Prompting 0.6985 + CoT Prompting 0.6873 + SGU-SQL 0.7395 CodeX + Few-shot Prompting 0.7167 + CoT Prompting 0.7198 + SGU-SQL 0.7418 ChatGPT + Few-shot Prompting 0.73947 + CoT Prompting 0.7552 + SGU-SQL 0.7846 GPT-4 + Few-shot Prompting 0.7665 + CoT Prompting 0.7834 + SGU-SQL 0.8791
 4.4 Model Analysis: Q3